Calculate the percent by mass of the solution containing 8.60 g of tha salt in 95.0 g of solution.

1
Expert's answer
2021-12-10T03:30:12-0500

%(m/m) = "\frac{mass of solute }{mass of solution}" × 100%


"\frac{8.60g}{95.0g}" × 100%


= 9.052 %